### v1

# How informative are summaries of the cosmic 21-cm signal?

11institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, 56125 Pisa, Italy
11email: david.prelogovic@sns.it
The cosmic 21-cm signal will bring data-driven advances to studies of the Cosmic Dawn (CD) and Epoch of Reionization (EoR). Radio telescopes such as the Square Kilometre Array (SKA) will eventually map the HI fluctuations over the first billion years – the majority of our observable Universe. With such large data volumes, it becomes increasingly important to develop ?optimal? summary statistics, allowing us to learn as much as possible about the CD and EoR. In this work we compare the astrophysical parameter constraining power of several 21-cm summary statistics, using the determinant of the Fisher information matrix, $\det\bm{F}$. Since we do not have an established ?fiducial? model for the astrophysics of the first galaxies, we compute for each summary the distribution of $\det\bm{F}$ across the prior volume. Using a large database of cosmic 21-cm lightcones that include realizations of telescope noise, we compare the following summaries: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; (v) an information-maximizing neural network (IMNN); and (vi) the combination of 2DPS and IMNN. Our best performing individual summary is the 2DPS, having relatively high Fisher information throughout parameter space. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. Our best results are achieved with the concatenation of the 2DPS and IMNN. The combination of only these two complimentary summaries reduces the recovered parameter variances on average by factors of $\sim$6.5 – 9.5, compared with using each summary independently. Finally, we point out that that the common assumption of a constant covariance matrix when doing Fisher forecasts using 21-cm summaries can significantly underestimate parameter constraints.
How informative are summaries of the cosmic 21-cm signal?
David Prelogović    Andrei Mesinger
(Received - - - / Accepted - - -)
Key Words.:
Methods: data analysis – Methods: statistical – Cosmology: theory – dark ages, reionization, first stars
1 Introduction
The cosmic 21-cm signal, corresponding to the spin flip transition of the ground state of HI, provides a window on the first billion years of the Universe’s evolution. This under-explored period witnessed fundamental cosmic milestones, including the Cosmic Dawn (CD) of the first galaxies and the final phase change of our Universe: the Epoch of Reionization (EoR). Current radio interferometers are putting increasingly tighter upper limits on its power spectrum (PS; e.g. Trott et al. 2020; Mertens et al. 2020; HERA Collaboration et al. 2023), while the upcoming Square Kilometre Array (SKA) [^1] https://www.skatelescope.org is promising to provide a 3D image of the EoR in the next decade(s) (e.g. Koopmans et al. 2015; Mesinger 2020). The unknown properties of the first galaxies and IGM structures are encoded in the timing and morphology of the cosmic 21-cm signal (e.g. Pritchard & Furlanetto 2007; McQuinn & O’Leary 2012; Visbal et al. 2012; Pacucci et al. 2014).
How do we infer these properties? The only statistically-robust way is through Bayesian inference: sampling from the theory and comparing forward models to the observation(s). However what should we use to compare the forward model to the data; i.e. how should we construct the likelihood? The cosmic 21-cm signal is intrinsically a 3D lightcone map. Performing inference directly on the lightcone would be incredibly challenging, due to the high-dimensionality of the data. Effectively, at each frequency the SKA should obtain a sky map comparable to current CMB maps, but with the full lightcone including thousands of such frequency slices (e.g. Loeb & Zaldarriaga 2004). [^2] Inference on high-dimensional lightcones is still possible in some cases, for example galaxy large scale structure (LSS) maps (e.g. Kitaura & Enßlin 2008; Jasche & Kitaura 2010; Jasche & Wandelt 2012, 2013; Leclercq et al. 2017; McAlpine et al. 2022; Dai & Seljak 2022; Bayer et al. 2022). Although such studies still compress the full galaxy observations, only treating them as biased tracers of the matter field, they are able to recover the phase information of our Universe’s initial conditions (so-called ?constrained realizations?). It remains to be seen if such studies can be extended to the Epoch of Reionization and Cosmic Dawn (though see Zhao et al. 2023b for a recent proof-of-concept study). Instead, the data is compressed into a summary statistic. Summary statistics usually involve some form of averaging, which increases the signal to noise (S/N) and can motivate assuming an approximately Gaussian likelihood when performing inference (e.g. Greig & Mesinger 2015; Gazagnes et al. 2021; Watkinson et al. 2022). [^3] The gaussianity of summaries constructed by averaging can be loosely motivated by appealing to the Central Limit Theorem. However, the theorem holds only for independent, identically distributed random variables. Unfortunately, we only have a single observable Universe; thus summaries of cosmological data sets (e.g. the power spectrum) often involve binning over correlated (i.e. not independent) and/or differently evolving fields (i.e. not identically distributed), resulting in non-Gaussian likelihoods (e.g. Prelogović & Mesinger 2023).
The most common choice of a 21-cm summary statistic is the spherically-averaged power spectrum (1D PS). The 1D PS maximizes the signal-to-noise (S/N) of an interferometric observation – a primary concern in the run up to a first detection (e.g. Greig & Mesinger 2018; Trott et al. 2020; Mertens et al. 2020; HERA Collaboration et al. 2023). Another common choice of a summary is the cylindrically-averaged (2D) PS, which includes additional information on the anisotropy of the cosmic 21-cm signal (e.g. Bharadwaj & Ali 2005; Barkana & Loeb 2006; Datta et al. 2012; Mao et al. 2012). Fourier space is also a natural basis to isolate the cosmic signal from instrumental effects and foregrounds, which primarily reside in a wedge region in the 2D PS (e.g. Morales et al. 2012; Vedantham et al. 2012; Trott et al. 2012; Parsons et al. 2014; Liu et al. 2014a, b; Murray & Trott 2018). However, the power spectrum ignores Fourier phases, which can encode significant information for a highly non-Gaussian signal such as 21-cm from the EoR/CD (e.g. Bharadwaj & Pandey 2005; Mellema et al. 2006; Shimabukuro et al. 2016; Majumdar et al. 2018; Watkinson et al. 2019). Because of this, several studies have also explored non-Gaussian statistics, such as the bispectrum (Shimabukuro et al., 2016; Majumdar et al., 2018; Watkinson et al., 2019, 2022), trispectrum (Cooray et al., 2008; Flöss et al., 2022), morphological spectra (Gazagnes et al., 2021), bubble size distribution (Lin et al., 2016; Giri et al., 2018; Shimabukuro et al., 2022; Doussot & Semelin, 2022; Lu et al., 2023), one-point statistics (Harker et al., 2009; Watkinson & Pritchard, 2015; Kittiwisit et al., 2022), genus topology (Hong et al., 2014), Wavelet Scattering Transform (Greig et al., 2022, 2023; Zhao et al., 2023a), Minkowski functionals and tensors (Gleser et al., 2006; Yoshiura et al., 2017; Chen et al., 2019; Kapahtia et al., 2019; Spina et al., 2021) and Betti numbers (Giri & Mellema, 2021; Kapahtia et al., 2021).
Is there an ?optimal? choice of summary statistic? One way to define ?optimal? is by how tightly we can recover cosmological/astrophysical parameters. This can be quantified through the Fisher information matrix (Fisher, 1935). The Fisher information is commonly used to make forecasts using pre-determined summary statistics, or even to define the summary statistic itself. Optimal summaries/algorithms that are based on the Fisher matrix include Karhunen-Loève methods and Massively Optimised Parameter Estimation and Data compression (MOPED) for linear transformations (e.g. Tegmark et al. 1997; Heavens et al. 2000), as well as non-linear generalizations like Information Maximizing Neural Networks (IMNNs, Charnock et al. 2018) and Fishnets (Makinen et al., 2023). However, these summary statistics come at the cost of interpretability, making the compression more of a ?black box? compared to physically-motivated and easy to interpret summaries like the PS. [^4] An interesting question is why would we even expect a truly optimal (lossless) compression to exist? The answer to this can be motivated through the manifold hypothesis (Fefferman et al., 2016), which states that physical probability distributions of the data often lie on a low-dimensional manifold. Optimal (lossless) compression would then ?extract? such manifolds from the original high-dimensional data. The above-mentioned algorithms do not attempt this directly, but aim at making compression as optimal as possible, given certain assumptions.
It is also important to note that, unlike $\Lambda$CDM, there is no obvious or unique choice to parametrize the astrophysics of early galaxies whose radiation determines the cosmic 21-cm signal. Neither do we have a good idea of a ?fiducial? model. Thus there is no guarantee that a summary statistic that maximizes Fisher information at a single parameter value for a single model parametrization would generalize well to other parameters / models.
Here we systematically compare the information content of several common 21-cm summaries on the basis of their Fisher information. These include: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; and (v) an IMNN. Several previous studies compared the constraining power of 21-cm summaries (e.g. Gazagnes et al. 2021; Watkinson et al. 2022; Greig et al. 2022; Zhao et al. 2023b; Hothi et al. 2023). Here we explore a broader range of summary statistics, introducing also IMNNs to the field. Unlike previous 21-cm Fisher studies, we do not make the simplifying assumption of a constant covariance matrix. Most importantly, we compare the summary statistics across a broad range of parameter space, instead of choosing only one or two fiducial values to make a mock 21-cm observation. This is very important to verify the robustness of the summary, since we do not have a good idea of a ?fiducial? astrophysical model for the first galaxies.
The paper is organized as follows. In section §2 we define the information content of a summary through the Fisher matrix and introduce the Fisher information distribution. In section §3 we introduce our 21-cm simulator and explain the summaries we consider in detail. In §4 we discuss our database of prior samples for evaluating the Fisher information. In §5 we present our results, quantifying the most informative summaries as well as the error introduced by the common assumption of a constant covariance matrix. Finally, we present concluding remarks and future prospects in §6. All quantities are quoted in co-moving units, and we assume a standard $\Lambda\mathrm{CDM}$ cosmology: $\left(\Omega_{\Lambda},\Omega_{\mathrm{M}},\Omega_{b},n,\sigma_{8},H_{0}\right)=\left(0.69,0.31,0.048,0.97,0.81,68\mathrm{\leavevmode\nobreak\ km}\mathrm{\leavevmode\nobreak\ s}^{-1}\mathrm{Mpc}^{-1}\right)$, consistent with the results from Planck Collaboration et al. (2020).
2 The Fisher matrix as a measure of information content
How do we measure the information content of a summary? If we have a theoretical model with a given parametrization, we can see how sensitive the summary is to changes in the astrophysical/cosmological parameters around some fiducial value. Specifically, we can calculate the Fisher matrix (e.g. Spall 2005):
$\bm{F}(\bm{\theta}^{*})_{mn}$ $=\int\mathrm{d}\bm{d}\,P(\bm{d}|\bm{\theta}^{*})\,\,\,\frac{\partial}{\partial\bm{\theta}_{m}}\ln P(\bm{d}|\bm{\theta}^{*})\cdot\frac{\partial}{\partial\bm{\theta}_{n}}\ln P(\bm{d}|\bm{\theta}^{*})$
$=\mathbb{E}\left[\left.\frac{\partial}{\partial\bm{\theta}_{m}}\ln P(\bm{d}|\bm{\theta})\cdot\frac{\partial}{\partial\bm{\theta}_{n}}\ln P(\bm{d}|\bm{\theta})\right|{\bm{\theta^{*}}}\right]\,,$ (1)
where $\bm{d}$ is the data summary (c.f. Fig. 1), $P(\bm{d}|\bm{\theta})$ is the likelihood, and $\mathbb{E}$ denotes the expectation value over the likelihood (i.e. an empirical average over many realizations) evaluated at a given point in parameter space $\bm{\theta}^{*}$. The Fisher information matrix provides the maximum constraining power, known as the Cramer-Rao bound (Rao, 1992; Cramér, 1999): $\operatorname{Var}(\hat{\bm{\theta}}_{m})\geq(\bm{F}^{-1})_{mm}$, where $\hat{\bm{\theta}}$ is an unbiased estimator of the parameters. In the multivariate case, one can equivalently write:
$\operatorname{Cov}(\hat{\bm{\theta}})\geq\bm{F}^{-1}\,,$ (2)
where the matrix inequality is interpreted as $\operatorname{Cov}(\hat{\bm{\theta}})-\bm{F}^{-1}$ being a positive semi-definite matrix. One can then prove the following inequality (see Appendix A):
$\det{\operatorname{Cov}(\hat{\bm{\theta}})}\geq\det{\bm{F}^{-1}}\,.$ (3)
Therefore, the determinant of the Fisher matrix measures the tightest parameter volume that can be constrained by the data. From now on, we will refer to $\det\bm{F}(\bm{\theta}^{*})$ as simply the Fisher information around $\bm{\theta}^{*}$ the higher its value, the more constraining the summary is around $\bm{\theta}^{*}$.
2.1 Gaussian approximation
In almost all cosmological scenarios, we do not explicitly know the full likelihood function $P(\bm{d}|\bm{\theta})$, nor its ?score? $\nabla_{\bm{\theta}}\ln P(\bm{d}|\bm{\theta})$. Instead for performing quick forecasts, it is common to approximate the likelihood as a Gaussian in data space $P(\bm{d}|\bm{\theta})\approx\mathcal{N}(\bm{d}|\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$, where the mean $\bm{\mu}(\bm{\theta})$ and covariance $\bm{\Sigma}(\bm{\theta})$ are general functions of the parameter space.
Assuming a Gaussian likelihood, one can compute the integral in Eq. 1 analytically (Appendix B; see also Tegmark et al. 1997; Vogeley & Szalay 1996):
$\bm{F}_{mn}$ $\equiv\bm{F}_{\bm{\mu},\,mn}+\bm{F}_{\bm{\Sigma},\,mn}$
$=\frac{\partial\bm{\mu}^{T}}{\partial\theta_{m}}\bm{\Sigma}^{-1}\frac{\partial\bm{\mu}}{\partial\theta_{n}}+\frac{1}{2}\operatorname{tr}\left[\bm{\Sigma}^{-1}\frac{\partial\bm{\Sigma}}{\partial\theta_{m}}\bm{\Sigma}^{-1}\frac{\partial\bm{\Sigma}}{\partial\theta_{n}}\right]\,.$ (4)
Here the first (second) term measures how the mean (covariance) of the data summary changes with respect to the parameters. Cosmological Fisher forecasts usually ignore one of the two terms (most often the second term $\bm{F}_{\bm{\Sigma}}$), effectively fixing either the mean or the covariance matrix to some fiducial value. However, we will see below that both terms are non-negligible for some summaries (e.g. Carron 2013).
2.2 Finite differencing
In order to numerically evaluate Eq. 4, we group the terms into ?fixed? ($\bm{\mu}$, $\bm{\Sigma}$, $\bm{\Sigma}^{-1}$) and ?differentiated? ($\nabla_{\bm{\theta}}\,\bm{\mu}$, $\nabla_{\bm{\theta}}\,\bm{\Sigma}$). We calculate ?fixed? terms from realizations of the summary around the same parameter value $\bm{\theta}^{*}$. To compute the mean, we can use the maximum likelihood estimator (MLE):
$\bm{\mu}(\bm{\theta}^{*})=\frac{1}{N}\sum_{i=1}^{N}\bm{d}(\bm{\theta}^{*})_{i}\,,$ (5)
where the index $i$ labels different realizations of initial conditions and other sources of stochasticity (e.g. telescope noise). It is known that such an estimator of the mean is unbiased. However for the covariance matrix the situation is more complex, as the MLE
$\bm{\Sigma}_{\text{MLE}}(\bm{\theta}^{*})=\frac{1}{N}\sum_{i=1}^{N}(\bm{d}(\bm{\theta}^{*})_{i}-\bm{\mu}(\bm{\theta^{*}}))(\bm{d}(\bm{\theta}^{*})_{i}-\bm{\mu}(\bm{\theta}^{*}))^{T}\,$ (6)
is biased. One can show that the estimators of the covariance and its inverse can be unbiased in the following way (e.g. Hartlap et al. 2007):
$\bm{\Sigma}(\bm{\theta}^{*})$ $=\frac{N}{N-1}\,\bm{\Sigma}_{\text{MLE}}(\bm{\theta}^{*})\,,$ (7)
$\bm{\Sigma}^{-1}(\bm{\theta}^{*})$ $=\frac{N-D-2}{N-1}\,\bm{\Sigma}_{\text{MLE}}^{-1}(\bm{\theta}^{*})\,,$ (8)
where $D$ is dimensionality of the summary $\bm{d}$. The correction prefactor for the inverse of the covariance is particularly important when $D$ is close to the number of samples $N$. Without it, both $\bm{F}_{\bm{\mu}}$ and $\bm{F}_{\bm{\Sigma}}$ would end up over-confident, with the effect on $\bm{F}_{\bm{\Sigma}}$ being quadratic (see Eq. 4).
A similar approach is used to estimate the ?differentiated parameters?. In the case in which the simulator is not differentiable, we can use finite differencing to estimate gradients of the mean and covariance:
$\frac{\delta\bm{d}(\bm{\theta}^{*})_{i}}{\delta\bm{\theta}_{m}}$ $\equiv\frac{\bm{d}(\bm{\theta}^{*}+\delta\bm{\theta}_{m}/2)_{i}-\bm{d}(\bm{\theta}^{*}-\delta\bm{\theta}_{m}/2)_{i}}{\delta\bm{\theta}_{m}}\,,$ (9)
$\left.\frac{\partial\bm{\mu}}{\partial\bm{\theta}_{m}}\right|_{\bm{\theta}^{*}}$ $=\frac{1}{M}\sum_{i=1}^{M}\,\frac{\delta\bm{d}(\bm{\theta}^{*})_{i}}{\delta\bm{\theta}_{m}}\,,$ (10)
$\left.\frac{\partial\bm{\Sigma}}{\partial\bm{\theta}_{m}}\right|_{\bm{\theta}^{*}}$ $=\frac{2}{L-1}\sum_{i=1}^{L}\left(\frac{\delta\bm{d}(\bm{\theta}^{*})_{i}}{\delta\bm{\theta}_{m}}-\left.\frac{\partial\bm{\mu}}{\partial\bm{\theta}_{m}}\right|_{\bm{\theta}^{*}}\right)\left(\bm{d}(\bm{\theta}^{*})_{i}-\bm{\mu}(\bm{\theta}^{*})\right)^{T}\,.$ (11)
Here, $\delta\bm{\theta}_{m}$ represents a small change in the parameter over which the derivative is computed and $\delta\bm{d}(\bm{\theta}^{*})_{i}/\delta\bm{\theta}_{m}$ corresponds to the numerical derivative of a given summary realization (i.e. fixing all sources of stochasticity).
2.3 The distribution of the Fisher information over parameter space
Below we will show examples of common 21-cm summaries evaluated at a ?fiducial? set of parameters, $\bm{\theta}_{\text{fid}}$. Unfortunately, the cosmic 21-cm signal does not yet have a well-defined fiducial model and our choice of $\bm{\theta}_{\text{fid}}$ is only a ?best guess?. Therefore in order to properly assess the quality of a particular summary, we sample many points $\bm{\theta}^{*}$ from the prior $P(\bm{\theta})$, calculating the Fisher information $\det\bm{F}(\bm{\theta}^{*})$ at each sample. We use this prior-weighted distribution of Fisher information as the main metric for assessing the quality of a given summary.
3 The cosmic 21-cm signal and its summary statistics
Figure 1: Schematic illustrating our pipeline for producing data summaries for a given astrophysical parameter combination (adapted from Prelogović & Mesinger 2023). Starting from a realization of the cosmological signal simulated with 21cmFAST, we remove the mean at each redshift and add a telescope noise realization corresponding to a 1000h observation with SKA1-Low. The resulting lightcone is compressed into different summaries. Explicitly-defined summaries are marked in blue, while neural network summaries are marked in red. This realization is computed at the fiducial parameter set, which used only for IMMN training and visualizations of the summaries. See text for more details.
For a given sample of astrophysical parameters, $\bm{\theta}^{*}$, we compute a realization of mock data using the following steps (illustrated in Fig. 1):
  1. 1.
    Sampling cosmological initial conditions, we simulate a realization of the 21-cm lightcone using the 21cmFAST code.
  2. 2.
    We subtract the mean from each frequency slice, mimicking an interferometric observation.
  3. 3.
    We simulate a $1000\mathrm{h}$ tracked scan with the SKA1-low telescope, including the corresponding $uv$ coverage and realization of the instrument noise.
  4. 4.
    We compress the resulting lightcone into a given summary observation.
We summarize these steps in more detail below.
The cosmic 21-cm signal is simulated using the semi-numerical code, 21cmFAST v3555https://github.com/21cmfast/21cmFAST (Mesinger & Furlanetto, 2007; Mesinger et al., 2011; Murray et al., 2020). In brief, we use the Park et al. (2019) galaxy model, varying five parameters that characterize the unknown UV and X-ray properties of high-$z$ galaxies through scaling relations:
  • $\bm{f_{*,10}}$ the fraction of galactic gas in stars, normalized at the halo mass of $10^{10}M_{\odot}$. The stellar to halo mass relation of the faint galaxies that drive reionization is well described by a power law: $M_{*}/M_{h}=f_{*,10}(M_{h}/10^{10}M_{\odot})^{0.5}(\Omega_{b}/\Omega_{m})$ (see Park et al. 2019 and references therein).
  • $\bm{f_{\text{esc},10}}$ the ionizing UV escape fraction, normalized at the halo mass $10^{10}M_{\odot}$. Again, the (mean) escape fraction is assumed to be a power law with halo mass, and here we fix its index to $\alpha_{\text{esc}}=-0.5$ (e.g. Qin et al. in prep).
  • $\bm{M_{\text{turn}}}$ the characteristic host halo mass below which galaxies are inefficient at forming stars, due to slow gas accretion, supernovae feedback, and/or photoheating.
  • $\bm{E_{0}}$ the characteristic X-ray energy below which photons are absorbed by the interstellar medium (ISM) of the host galaxy.
  • $\bm{L}_{\bm{X}<2\mathrm{keV}}/\mathrm{SFR}$ the X-ray soft band (in the energy range $E_{0}$ – 2 keV) luminosity per star formation rate (SFR), in units of $\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}$. The SFR is taken to be: $\dot{M}_{*}=M_{*}/(0.5H(z)^{-1})$, where the Hubble time, $H(z)^{-1}$, also scales with the dynamical time at the virial radius.
Our fiducial, $\bm{\theta}_{\text{fid}}$ parameter vector corresponds to the following: $\log_{10}f_{*,10}=-1.3$, $\log_{10}f_{\text{esc}}=-1$, $\log_{10}M_{\text{turn}}[M_{\odot}]=8.7$, $\log_{10}L_{X<2\mathrm{keV}}/\text{SFR}[\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}]=40$, $E_{0}[\mathrm{keV}]=0.5$. This particular choice is only used when training the IMMN, and for the visualizations of the summaries below. For more details on the model and its motivations, see Park et al. 2019 and references therein.
To simulate the observation we follow the procedure described in Prelogović & Mesinger (2023) (c.f. Fig. 1). In brief, we calculate the 21-cm signal from redshifts 30 to 5, on a 300 Mpc coeval cube, with 1.5 Mpc resolution. The snapshots of the evolved coeval cube are then stacked to produce the final lightcone of the signal. After the subtraction of the mean for every sky-plane slice, we add thermal noise corresponding to a a $6\,\mathrm{h/day}$, $10\,\mathrm{s}$ integration time tracked scan with SKA1-Low for a $1000\,\mathrm{h}$ observation in total. Before computing each summary, we additionally smooth the lightone with a box-car filter obtaining a final resolution of 6 Mpc. This final step does not impact our results as it is comparable to the SKA1-Low beam, but is needed in order to minimize GPU memory usage for certain summaries (see below for more details). We also note that in this work we make the optimistic assumption of perfect foreground removal, thus exploring the maximum future potential of our set of summary statistics.
The resulting 3D lightcone is compressed into a given summary statistic. We explore the following summaries in this work:
  • 1DPS
    - spherically-averaged 1D power spectrum,
  • 2DPS
    - cylindrically-averaged 2D power spectrum,
  • Wavelets
    - 2D wavelet scattering transform,
  • RNN
    - recurrent neural network,
  • IMNN
    - information maximizing neural network.
  • 2DPS + IMNN
    - concatenation of the 2DPS and IMNN summaries.
The first four are explicitly defined while the following two are ?learned? by neural networks. The last, 2DPS + IMNN, is a combination of the two individual summaries. We describe them in detail below.
3.1 Spherically-averaged (1D) Power Spectrum
Figure 2: 1D power spectrum for the fiducial lightcone shown in Figure 1. Different colors correspond to different wavemodes while solid (dashed) correspond to with (without) noise. Comparing the solid and dashed lines, we see that the signal is noise dominated at higher $k$ modes and redshifts (see also Fig. 1). All power spectra are computed from comoving cubes extracted from the lightcone, centered on the redshifts indicated by the vertical dashed lines.
As discussed in the introduction, the spherically-averaged (1D) power spectrum, 1DPS, is the most common choice of a summary statistic. Despite the cosmic 21-cm signal being non-Gaussian, the 1DPS is a natural choice when seeking to maximize the S/N of an interferometric first detection.
The 1DPS of the 21-cm signal $\delta T_{b}(\bm{x},z)$, where $\bm{x}$ is the sky-plane coordinate and $z$ the redshift, is defined as:
$\delta\bar{T}^{2}_{b}\Delta_{21}^{2}(k,z^{\prime})\equiv\frac{k^{3}}{2\pi^{2}V}\left\langle\left|\delta T_{b}(\bm{k},z^{\prime})-\delta\bar{T}_{b}(z^{\prime})\right|^{2}\right\rangle_{k}\,.$ (12)
[…]
3.2 Cylindrical (2D) Power Spectrum
Figure 3: 2D power spectrum of the fiducial model shown in Figure 1. Different columns show the 2DPS for the pure cosmic signal and including thermal noise, as labeled on the top. Each pair corresponds to a different central redshift, coinciding with those used in the 1DPS summary (see Fig. 2). Furthermore, the bins between 1DPS and 2DPS coincide as close as possible. As in the 1DPS case, one can notice that high-$k$ modes and high redshifts are noise dominated.
The cylindrically-averaged 2D PS distinguishes between sky-plane ($k_{\perp}$) and line-of-sight ($k_{\|}$) modes. Redshift space distortions (e.g. Bharadwaj & Ali 2005; Barkana & Loeb 2005) and lightcone evolution (e.g. Greig & Mesinger 2018; Mondal et al. 2018) result in an anisotropic cosmic signal. Therefore, the 2DPS should encode more physical information compared to the 1DPS. Moreover, the instrument/foregrounds are better characterized in the 2D PS (e.g. see the review in Liu et al. 2014b), making it a natural summary observable for preliminary, low S/N measurements.
The 2DPS is defined as:
$\delta\bar{T}^{2}_{b}\Delta_{21}^{2}(k_{\perp},k_{\|},z^{\prime})\equiv\frac{k^{2}_{\perp}\,k_{\|}}{4\pi^{2}V}\left\langle\left|\delta T_{b}(\bm{k},z^{\prime})-\delta\bar{T}_{b}(z^{\prime})\right|^{2}\right\rangle_{k_{\perp},k_{\|}}\,,$ (13)
where the Fourier transform is performed on the same lightcone chunks as in the 1D case, with the difference being that the expectation value is calculated over $(k_{\perp},k_{\|})$ bins. Here we pick $3\times 3$ log-spaced bins, aligned to the 1D case as close as possible. The 2DPS for the fiducial lightcone (see Fig. 1) is shown in Fig. 3. The panels correspond to the mean redshifts of the lightcone chunks, coinciding with the ones of the 1DPS (see Fig. 2). The columns correspond to the cases with and without noise, as labeled at the top. We can verify that both the cosmic signal and the noise are anisotropic. For the cosmic signal, this is most evident for $z\lesssim 10$ where the noise is subdominant. For the noise, the same can be clearly seen for higher redshifts, where the noise dominates the signal at high $k_{\|}$ much more than in high $k_{\perp}$.
3.3 Wavelet Scattering Transform
The wavelet transform (Gabor, 1946; Goupillaud et al., 1984; Trott et al., 2012) and wavelet scattering transform (WST) (Mallat, 2011) are based on convolutional filters designed on a Fourier basis with an additional Gaussian envelope along the frequency direction. As such, they capture local features in both image and Fourier domains, which is often relevant in audio and image processing. Higher-order convolutions combined with neural networks (wavelet-based CNNs) have been very successful for such purposes (Bruna & Mallat, 2012; Sifre & Mallat, 2013; Anden & Mallat, 2014). The WST has been extensively explored in cosmological (e.g. Cheng et al. 2020) and 21-cm (e.g. Greig et al. 2022, 2023; Zhao et al. 2023a; Hothi et al. 2023) analyses.
Here we briefly introduce how the WST is defined. It represents a convolution of an input image $I(\bm{x})$ by a set of wavelet (here specifically Morlet) filters. They are characterized by the physical scale $j$ and rotation moment $l$. The actual physical scale of the filter $\psi^{j,l}$ will then be $2^{j}$ pixels, and its orientation angle $l\cdot\pi/L$. The wavelet coefficients correspond to the average of the convolved images, to which we refer as simply wavelets. Note that in this work we restrict ourselves to the 2D WST (see e.g. Zhao et al. 2023a for recent work using the 3D transform).
The first and second order wavelets can be written as:
$s_{1}^{j_{1},l_{1}}$ $=\langle|I(\bm{x})*\psi^{j_{1},l_{1}}|\rangle\,,$ (14)
$s_{2}^{j_{1},l_{1},j_{2},l_{2}}$ $=\langle||I(\bm{x})*\psi^{j_{1},l_{1}}|*\psi^{j_{2},l_{2}}|\rangle\,.$ (15)
One can see the similarity between first order wavelets and the PS by writing $\mathrm{PS}(\bm{k})=\langle|I(\bm{x})*\psi^{\prime}|^{2}\rangle$, with $\psi^{\prime}=e^{-i\bm{k}\cdot\bm{x}}$. To extract only cylindrically-averaged information imprinted in the 21-cm signal (analogous to calculating 2DPS in a Fourier basis), one can average coefficients over the rotations:
$S_{1}^{j_{1}}$ $=\langle s_{1}^{j_{1},l_{1}}\rangle_{l_{1}}\,,$ (16)
$S_{2}^{j_{2}}$ $=\langle s_{2}^{j_{1},l_{1},j_{2},l_{2}}\rangle_{l_{1},l_{2}}\,.$ (17)
For further information about the WST and wavelet coefficients we refer the reader to Cheng et al. (2020); Greig et al. (2022, 2023).
Our setup for calculating wavelets is the following. We first select a series of 8 images, centered at the same redshifts as for the power spectra. For each image we calculate the first and second-order wavelets and combine them for the final summary. As the WST scale $2^{J}\times 2^{J}$ has to be smaller than the sky-plane images ($50\times 50$), we set $J=4$. Further, we fix $L=4$ as a common choice (Cheng et al., 2020; Greig et al., 2022). For computing convolutions, we use the scattering_transform [^6] https://github.com/SihaoCheng/scattering_transform package (Cheng et al., 2020). For each redshift slice, we average the first and second order coefficients computed on a rolling 32x32 window. Our final WST summary then consists of $J=4$ $S_{1}$ and $J(J+1)/2=10$ $S_{2}$ coefficients for each of the 8 redshift slices, for a total of 112 numbers.
3.4 Recurrent Neural Network
A Recurrent Neural Network (RNN) is a type of Neural Network (NN) architecture designed to efficiently encode sequential or time-evolving information. In our previous work (Prelogović et al., 2022), we used RNNs along the redshift axis in combination with Convolutional NNs (CNNs) along the sky-plane, to capture the anisotropic evolution of the cosmic 21-cm signal along the lightcone. Specifically, we used a long short-term memory (LSTM, Hochreiter & Schmidhuber (1997); Shi et al. (2015)) RNN, which is more stable to train compared with older RNN versions (see e.g. the review in Schmidt 2019). Using RNNs resulted in a more accurate regression of astrophysical parameters, compared to only using CNNs (e.g. Gillet et al. 2019; La Plante & Ntampaka 2019; Mangena et al. 2020; Kwon et al. 2020; Zhao et al. 2022a; Heneka 2023).
In this work, we use the flagship model from Prelogović et al. (2022), SummaryRNN, together with the exact weights from that work. The model has been trained in a supervised manner as a regressor. The database used for training had the same telescope simulator, but used an older version of the 21cmFAST simulator with a different astrophysical parametrization.
Our RNN summary in this work is the 32-dim output from the first dense layer of the network (see Table B3 in Prelogović et al. 2022, where convolutional and recurrent layers are followed by dense layers, ending with prediction of four parameters). Using such a higher-dimensional layer instead of the final parameter regressor output (e.g. Zhao et al. 2022a), makes the summary more sensitive to the general lightcone features instead of the specific model and/or parametrization used to create the training set. This is similar to so-called ?transfer learning? (e.g. Jiang et al. 2022), where one uses a NN that has been pre-trained on a different domain (i.e. database) and re-trains its last layers on a new domain. As our original RNN takes as input $25\times 25$ sky-plane slices, we roll a $25\times 25$ window over each redshift slice, averaging the corresponding dense layer outputs to obtain the final RNN summary.
3.5 Information Maximizing Neural Network
An Information Maximizing Neural Network (IMNN; Charnock et al. 2018) [^7] https://github.com/tomcharnock/IMNN is an unsupervised learning algorithm, where the NN is specifically trained to maximize the Fisher information $\det\bm{F}$ of the summary compression, at a single fixed point in parameter space $\bm{\theta}_{\text{fid}}$. IMNNs have been previously applied to galaxy LSS maps (e.g. Makinen et al. 2022) and catalog-free modeling of galaxy types (Livet et al., 2021). Here we apply them for the first time to 21-cm maps of the EoR and CD. In what follows, we outline the main ingredients of the algorithm and the training procedure.
Taking some NN architecture (in our case a simple CNN), one compresses the lightcone $\bm{l}$ into a summary vector $\bm{d}$
$\bm{d}=\mathrm{NN}(\bm{l})\,.$ (18)
The dimensionality of the summary is equal to the parameter space dimensionality ($D$). This is motivated by the fact that, for a given model parametrization, the ?score? of the real likelihood function $\nabla_{\bm{\theta}}\ln P(\bm{l}|\bm{\theta})$ is sufficient for optimal parameter recovery, and thus the dimensionality of the summary which maximizes information around one fiducial point can be the same as the dimensionality of the parameter space (e.g. Heavens et al. 2000; Charnock et al. 2018).
We train the IMMN using the following loss function (c.f. Charnock et al. 2018; Makinen et al. 2021):
$\bm{F}_{\text{IMNN}}(\bm{\theta}_{\text{fid}})_{mn}$ $\equiv\left.\frac{\partial\bm{\mu}^{T}}{\partial\bm{\theta}_{m}}\right|_{\bm{\theta}_{\text{fid}}}\,\left.\frac{\partial\bm{\mu}}{\partial\bm{\theta}_{n}}\right|_{\bm{\theta}_{\text{fid}}}$ (19)
$L_{\mathcal{F}}({\bm{\Sigma}(\bm{\theta}_{\text{fid}})})$ $\equiv\|\bm{\Sigma}(\bm{\theta}_{\text{fid}})-\mathbb{1}\|_{\mathcal{F}}$ (20)
$L_{2}(W_{\text{NN}})$ $\equiv\|W_{\text{NN}}\|^{2}$ (21)
$\mathcal{L}$ $=-\ln\det\bm{F}_{\text{IMNN}}$
$+\lambda\,L_{\mathcal{F}}({\bm{\Sigma}})\tanh L_{\mathcal{F}}(\bm{\Sigma})$ (22)
$+\lambda_{W}\,L_{2}(W_{\text{NN}})$
where
$\bm{\mu}(\bm{\theta_{\text{fid}}})$ $=\frac{1}{N}\sum_{i=1}^{M}\bm{d}(\bm{\theta}_{\text{fid}})_{i}\,,$ (23)
$\bm{\Sigma}(\bm{\theta}_{\text{fid}})$ $=\frac{1}{N-1}\sum_{i=1}^{N}\left(\bm{d}(\bm{\theta}_{\text{fid}})_{i}-\bm{\mu}(\bm{\theta_{\text{fid}}})\right)\left(\bm{d}(\bm{\theta}_{\text{fid}})_{i}-\bm{\mu}(\bm{\theta}_{\text{fid}})\right)^{T},$ (24)
$\left.\frac{\partial\bm{\mu}}{\partial\bm{\theta}_{m}}\right|_{\bm{\theta}_{\text{fid}}}$ $=\frac{1}{M}\sum_{i=1}^{M}\,\frac{\bm{d}(\bm{\theta}_{\text{fid}}+\delta\bm{\theta}_{m}/2)_{i}-\bm{d}(\bm{\theta}_{\text{fid}}-\delta\bm{\theta}_{m}/2)_{i}}{\delta\bm{\theta}_{m}}\,,$ (25)
$L_{\mathcal{F}}({\bm{\Sigma}(\bm{\theta}_{\text{fid}})})$ is the Frobenius norm between the covariance and the unit matrix $\bm{\mathbb{1}}$, $L_{2}(W_{NN})$ is the $l_{2}$ norm of the networks’ weights. The final loss is then the sum of the negative logarithm of the Fisher information, together with the covariance and weights regularizers. The $\tanh$ term in the covariance regularizer is used to turn-off the regularization once the covariance becomes close to the unity. The main goal of the IMNN is therefore to maximize the Fisher information, while keeping the covariance of the summary fixed to unity. [^8] For any summary $\bm{d}$ of dimensionality D, one can define a new summary, $\bm{d}^{\prime}=\bm{a}+\bm{B}\bm{d}$, where $\bm{a}$ is a constant D-dim vector serving to normalize $\bm{d}^{\prime}$ to be zero mean, and $\bm{B}$ is a non-singular $D\times D$ matrix, serving to normalize $\bm{d}^{\prime}$ to be unit variance. One can then show that $\ln P(\bm{d}|\bm{\theta})=\ln P(\bm{d}^{\prime}|\bm{\theta})+\ln\det\bm{B}\,$, or equivalently $\partial\ln P(\bm{d}|\bm{\theta})/\partial\bm{\theta}_{m}=\partial\ln P(\bm{d}^{\prime}|\bm{\theta})/\partial\bm{\theta}_{m}$. This means that one always has the freedom to transform the summary $\bm{d}$ to be zero-mean, unit-variance at a single point $\bm{\theta_{\rm fid}}$, without changing the Fisher information anywhere in the parameter space. Note that the Fisher matrix is calculated only using the first term and without covariance inverse between the two derivatives (see Eq. 4). Both of these simplifications were needed in order to stabilize the training of the IMMN, and they are correct as long as the covariance is kept at unity. [^9] The original work by Charnock et al. (2018) had slightly different loss function, by having the inverse of the covariance in the Fisher term and the additional Frobenius norm between $\bm{\Sigma}^{-1}$ and $\bm{\mathbb{1}}$. We found that these choices destabilized the training and are not crucial for the convergence (see also Makinen et al. 2021).
For the IMNN training, we create a database around the fiducial $\bm{\theta}_{\text{fid}}$ (see Figs. 1 and 4) for $N=M=1024$ (Eqs. 23, 24, 25), with $11264$ realizations in total. The training and validation sets are formed by splitting the database in half, following the original work of Charnock et al. (2018). We train the network using stochastic gradient descent with a batch size $b=128$, regularization strengths $\lambda=10$, $\lambda_{W}=10^{-9}$ and the Adam optimizer. To increase the effective database size and improve the training, the telescope noise is added to the simulations ?on-the-fly?. The final model has been trained on $8$ V-100 GPUs in parallel, for $\approx 120\mathrm{k}$ epochs and $\approx 20\mathrm{k}$ GPUh. The IMNN code developed for the purposes of this work we make publicly available under 21cmIMNN [^10] https://github.com/dprelogo/21cmIMNN package.
3.6 2DPS + IMNN
Finally, we also compute the Fisher information of the concatenation of the individual 2DPS and IMNN summaries. As we shall see in the following section, these are the two best performing individual summaries. Note that one can cleanly combine summaries when performing Simulation Based Inference (SBI), as the likelihood does not need to be specified explicitly. The main restriction is the size of the required database and the stability of the SBI training (e.g. Cranmer et al. 2020; Lueckmann et al. 2021). When computing the Fisher information in this work, we are restricted to these two summaries due to the limits imposed by our database size. In particular, the minimal number of samples needed to estimate the covariance matrix is proportional to the dimensionality of the total summary (e.g. Hartlap et al. 2007). Furthermore, the dimensions of the final summary should be uncorrelated, as otherwise the covariance is singular and non-invertible. This can result in incorrect Fisher matrix estimation (see however Tegmark et al. 1997 for possible alternative solutions).
4 Database
Figure 4: Prior volume of the astrophysical parameter space. Dashed lines demarcate the fiducial parameter values to train the IMNN, 2D contours enclose 95% of the prior volume, and the points denote the 152 prior samples at which we compute the Fisher information for our summaries. Note that all units and $\log_{10}$ factors have been removed for clarity - see the main text and Fig. 1 for reference on the units.
Since we do not have a strongly-motivated fiducial choice for our galaxy parameters, we want to compare the Fisher information of 21-cm summaries across our prior volume. In order to do so, we construct a ?Fisher database?, sampling 152 parameter combinations from our prior. We use a flat prior for the X-ray emission parameters, $L_{X}$ and $E_{0}$, over the ranges shown in Fig. 4 (for more motivation of these choices see Fragos et al. 2013; Lehmer et al. 2016; Das et al. 2017). The remaining parameters characterizing star formation and ionizing emission of galaxies are constrained by observed UV luminosity functions (e.g. Bouwens et al. 2015, 2017; Oesch et al. 2018) as well as estimates of the reionization history (e.g. McGreer et al. 2015; Planck Collaboration et al. 2016). As our model is nested in the 6D Park et al. (2019) model, we use the posterior from that work, calculating the corresponding conditional distributions over $f_{*,10}$, $f_{\text{esc},10}$ and $M_{\text{turn}}$, fixing the remaining parameters to their fiducial values, as discussed in the previous section. To compute the conditional distributions, we use the conditional_kde [^11] https://github.com/dprelogo/conditional_kde code.
Our prior and samples are shown in Fig. 4. The diagonal panels illustrate the 1D prior PDFs, while the contours in the lower left enclose 95% of the 2D distributions. The 152 samples are shown as points, while the dashed vertical and horizontal lines demarcate the fiducial choice used to train our IMNN.
We simulate $N=130$ realizations for each of the 152 parameter samples, $\bm{\theta^{*}}$, in order to estimate the mean and the covariance of each summary (Eqs. 5, 6), plus an additional $1502\times 5\times M=$ 150 realizations at $\bm{\theta^{*}}\pm\delta\bm{\theta}_{m}/2$ to estimate the derivative of the mean (Eq. 10) along each of the 5 parameter basis vectors. We note that no new simulations are needed to estimate the derivative of the covariance matrix, since we use the same $L=M=15$ realizations that are used to compute the first Fisher term mentioned above (c.f. Eq. 11 and associated discussion).
5 Results
We now present the main results of this work. We first discuss the training of the IMNN in Section 5.1, then compare the performance of all of our 21-cm summaries in Section 5.2, adding a final discussion on the importance of including the second Fisher term (c.f. Eq. 4) in Section 5.3. Throughout, our figure of merit will be the PDF of the Fisher information ($\log_{10}(\det\bm{F})$), over the prior volume (see previous section). We quote both the median and the variance of this distribution. A high value of the median means that the typical parameter value is well constrained by the summary, while a narrow variance means that the constraining power of the summary is relatively constant over the prior volume. We note that the Fisher information is a measure of the average constraining power over our five parameters; we do not discuss the relative constraints on individual parameters in this work.
5.1 IMNN training
Figure 5: Fisher information of the IMNN during training. Dark (light) blue curves show the Fisher information for the training (validation) set as the IMNN is trained at $\bm{\theta}_{\rm fid}$. The violin plots show the distribution of the Fisher information over the full prior volume excluding $\bm{\theta}_{\rm fid}$, $P(\theta)$, at every interval of 10k training epochs. Black dots and error bars inside each violin denote the median and 68% C.L. of the distributions. The divergence between the Fisher information computed at $\bm{\theta}_{\rm fid}$ and its PDF over the remaining prior volume clearly indicates that the IMNN is over-specializing to the specific parameter combination, $\bm{\theta}_{\rm fid}$, used in the training.
In Fig. 5 we show the training results for the IMNN. Blue lines show the training and validation Fisher information (see the first term of the IMNN loss, Eq. 22) at the fiducial point $\bm{\theta}_{\text{fid}}$. We recall that the training and validation sets each consist of 5632 lightcone realizations, evaluated at $\bm{\theta}_{\text{fid}}$ with different cosmic initial conditions and telescope noise seeds. The Fisher information of the training and validation sets are largely overlapping meaning the network does not overfit to the training set. We see that by the end of the training, the IMNN has learned to improve the Fisher information by 40(!) orders of magnitude, corresponding to parameter constraints that are improved by an average factor of $\sim 10^{8}$ in the variance!
Is the compression learned by the IMNN comparably informative at other parameter combinations? We can answer this question by comparing the solid lines of Fig. 5 to the violin distributions. The latter show the distribution of the Fisher information over the full prior volume, $P(\theta)$, at every interval of 10k training epochs, excluding $\bm{\theta}_{\rm fid}$. We see that although the IMNN is not overfitting in the classical sense (the Fisher information of the validation set keeps increasing), it quickly ?over-specializes? to the specific parameter value used for training, $\bm{\theta}_{\rm fid}$. Thus the compression learned by the IMNN does not generalize well across parameter space. [^12] One solution could be to train the IMNN around several different points. However, this is extremely expensive in our application, both in terms of database size and GPU time. Furthermore, it is not clear if such procedure would result in a meaningful summary, as the concept of one point training is inherently built into the IMNN. Alternatively, it might be possible to construct a generalized IMNN, maximizing the Fisher information throughout parameter space; however, it is not obvious how this could be done. The PDF of the Fisher information over the prior volume does not change significantly after epoch 40k, with the median beginning to decrease beyond epoch 90k. We take epoch 90k as the final state of the IMNN in the remainder of this work.
5.2 Comparison of all summaries
Figure 6: Distribution of the Fisher information ($\det\bm{F}$) across the prior volume, for all summaries considered in this work. For the IMNN summary, we use the weights that give the highest median value (corresponding to 90k epochs; see Fig. 5). On the top axis we denote the corresponding relative improvement in the single parameter variance (Eq. 26), normalized to its median value of the 2DPS summary. The 2DPS has the largest median value of all of the individual summaries, while the combination of the 2DPS + IMNN dramatically outperforms every individual summary. See Table 1 for quantitative values.
summary $\text{median}^{+34\%}_{-34\%}$ $\sigma^{2}/\sigma^{2}_{\text{2DPS}}$
RNN $\phantom{0}9.0_{-2.1}^{+1.6}$ 4.17
Wavelets $\phantom{0}9.6_{-3.1}^{+2.9}$ 3.16
IMNN 90k $11.8_{-5.7}^{+8.2}$ 1.15
1DPS $11.8_{-3.4}^{+3.1}$ 1.15
2DPS $12.1_{-3.3}^{+3.1}$ 1.00
IMNN 90k + 2DPS $15.9_{-4.9}^{+5.2}$ 0.17
Table 1: Companion table to Fig. 6, summarizing the Fisher information probability distribution across the prior volume for each summary. In particular, the first column represents values of $\log_{10}(\det\bm{F})$ for the median and $68\%$ of probability volume around it. The second column shows the single parameter variance (Eq. 26), which serves as a guide to how much would the recovered variance of a parameter increase or decrease if one uses that particular summary instead of the 2DPS.
In Fig. 6 we show the main result of this work: the distribution of Fisher information across the prior volume, for all of the summary statistics we consider. On the top axis we denote the corresponding relative improvement in constraints on the single parameter variance $\sigma^{2}$. We define it as an effective variance in each dimension with no correlation, giving the same Fisher information. For a 5-dim parameter space we have:
$\sigma^{2}=(\det\bm{F})^{-1/5}\,.$ (26)
Here, one can ignore the difference in the units (i.e. scaling) between parameter dimensions, see Appendix C for more details. Specifically, the top axis of Fig. 6 denotes the single parameter variance normalized to its median value of the 2DPS summary. To give an example, a factor of $10$ times larger Fisher information for a 5-dim parameter space would result in a $37\%$ smaller single parameter variance. The median values and 68% C.L. of the Fisher information and single parameter variance are also listed in Table 1.
Among the individual summaries, the 2DPS is a clear winner. In comparison, the 1DPS has roughly a factor of two smaller Fisher information across the prior volume, which translates to a $\sim 15\%$ larger single parameter variance (Eq. 26). This shows that indeed the anisotropy of the cosmic 21-cm signal helps in parameter recovery, though not very significantly. The distribution of the 2DPS Fisher information is also relatively narrow, indicating its constraining power does not vary enormously across parameter space, compared to most other summaries.
The median Fisher information of the IMNN is comparable with that of the 1DPS. However, the IMNN results in the widest distribution of all of the summaries. This means that the compression that the IMNN learned at $\bm{\theta}_{\rm fid}$ can be much more informative but also much less informative, depending on the parameter choice. Therefore, if one has a good idea of where the maximum likelihood value will be (e.g. from complementary observations; Park et al. 2019; Abdurashidova et al. 2022; HERA Collaboration et al. 2023; Breitman et al. 2024), it could be beneficial to train the IMNN at that ?best guess? value and use the resulting summary in inference. An optimal strategy could be to perform inference using the 2DPS in order to obtain the maximum likelihood, train the IMNN at this ML parameter combination, then perform inference again using the resulting IMNN summary.
The wavelet summary results in a median Fisher information that is a factor of $\sim$300 smaller than that of the 2DPS at their median values (factor of $\sim$3 increase in the single parameter variance (Eq. 26). Intuitively, one would expect more information content in wavelets compared with the PS, as they effectively include higher order correlations. However, our WST followed the definition in Greig et al. (2022, 2023), in which the wavelets are only computed at one slice for each redshift chunk. Thus, the physics-rich, line-of-sight modes inside each redshift bin are lost for this choice of WST. The weaker performance of the WST compared to the power spectra implies that there is more information in the line-of-sight modes than there is in higher-order correlations of the transverse modes.
We note that recent works also using 2D wavelets achieved better results than we do here (Greig et al., 2022; Hothi et al., 2023). However, their analysis differs from ours. In particular, they use more redshift bins and they do not average the coefficients over the sky-plane (rolling the filter as described above). We confirm that without the additional smoothing that we perform in this work, the Gaussian approximation for the likelihood that is intrinsic to the Fisher estimate (c.f. Eq. 4) is notably worse, for both the wavelets and the RNN. Therefore, the higher wavelet information found by Greig et al. (2022); Hothi et al. (2023) could be partially due to a less appropriate application of the Fisher estimate. Alternatively, the information content could be improved by using 3D wavelets instead of 2D ones (see e.g. Zhao et al. 2023a).
Of the individual summaries, the RNN performs the worst with a median Fisher information that is three orders of magnitude smaller than that of the 2DPS. However, its distribution is the most homogeneous across the prior volume (i.e. the narrowest). This could be interpreted as sacrificing the overall information content for robustness throughout the parameter space. Indeed the RNN was trained as a regressor over a large parameter space, and so it is understandable that it is much more consistent with respect to the IMNN which was trained on a single parameter combination. The overall low median information content of the RNN compression could be due to the fact that it was trained on a different parametrization of the cosmological signal than used in this work. We expect that retraining the RNN as a regressor on the same parametrization would improve the median information content. Indeed, Zhao et al. (2022a) find that a regressor CNN slightly outperforms the 1DPS, resulting in tighter parameter constraints when using two different parameter combinations for the mock observation. However, it is likely that this increase in the information content would come at the cost of robustness (i.e. a broader PDF), as the RNN over-specializes to features inherent in that model’s parametrization and its training set.
Finally, our concatenation of the 2DPS and IMNN significantly outperforms all of the individual summaries (c.f. the red curve in Fig. 6). As we saw above, these two summary statistics are highly complimentary: the 2DPS is robust (i.e. a narrow information PDF), while the IMNN summary can be much more constraining for some parameters but much less for others (i.e. it has a very broad information PDF). The distribution of Fisher information of 2DPS + IMNN retains the narrowness of the 2DPS, but is shifted to the right, including the high-value tail from the IMNN.  [^13] We note that in the case in which some value of the IMNN summary would be completely degenerate with some 2DPS bin, the resulting covariance matrix would be singular and therefore non-invertible. This would make estimating the Fisher matrix unreliable. However, as the IMNN compression and 2DPS are very different summaries, this is unlikely to be a problem. The combined summary is a factor of $\sim$6.5–9.5 better in the single parameter variance (Eq. 26), averaged over the prior samples. Combining multiple 21-cm statistics has already been proven to tighten parameter constraints (e.g. Gazagnes et al. 2021; Watkinson et al. 2022). However, previous work using traditional inference was limited by the need to derive a tractable likelihood which becomes impossible for non-trivial summaries. Fortunately, recent advances in applying SBI to 21-cm (e.g. Zhao et al. 2022a, b; Prelogović & Mesinger 2023; Saxena et al. 2023) will allow us to cleanly combine multiple summary statistics without having to explicitly define the likelihood. Unfortunately, the neural density estimation that is intrinsic to SBI can be unstable and/or require large training sets, if the summary statistics become very high dimensional. This work illustrates how combining only two complimentary statistics can result in much tighter, more robust parameter constraints than would be available using either individually.
5.3 One-term vs. full Fisher matrix
Figure 7: Distributions of Fisher information for the 1DPS (red) and 2DPS (blue). Solid lines account for both terms in the Fisher matrix (c.f. Eq. 4), while the dashed lines correspond to the common simplification of a constant covariance (i.e. including only the first term in Eq. 4). The assumption of a constant covariance underpredicts the Fisher information by a factor of $\sim 2$. For a 5-parameter model, this translates to a $\sim 15\%$ larger single parameter variance (Eq. 26).
Figure 8: Analogous to Fig. 7, but for the other 21-cm summary statistics.
The full Fisher matrix of a Gaussian likelihood can be separated in two terms $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ (see Eq. 4), where the first term includes the gradient of the mean and the second term the gradient of the covariance matrix (e.g. Appendix B; Tegmark et al. 1997). All of our previous results were calculated using the full expression, including both terms. However, it is common in the literature to ignore one of the terms, by approximating the mean or the covariance matrix as a constant (e.g. Carron 2013). In most of cosmological Fisher forecasts using the power spectrum (e.g. Sailer et al. 2021; Abazajian et al. 2022; Euclid Collaboration et al. 2022; Mason et al. 2023; Bykov et al. 2023), the second term, $\bm{F}_{\bm{\Sigma}}$, is neglected. [^14] For examples of ignoring the first term, see e.g. Abramo (2012); d’Assignies D et al. (2023). In some cases, the covariance matrix can indeed be independent or only weakly dependent on the parameters, justifying ignoring the second Fisher term. However, the importance of the $\bm{F}_{\bm{\Sigma}}$ term for 21-cm Fisher forecasts has not yet been investigated (e.g. Greig et al. 2022; Balu et al. 2023; Mason et al. 2023; Hothi et al. 2023). Indeed in Prelogović & Mesinger (2023) we showed that assuming a parameter-independent covariance matrix of the 21-cm 1DPS likelihood can bias the posterior estimation. In this section, we quantify the importance of the $\bm{F}_{\bm{\Sigma}}$ term in determining the Fisher information [^15] Assuming a constant covariance matrix can result in other biases at the level of the full Fisher matrix. For example, it could impact parameter degeneracies and/or improve constraints of some parameters at the cost of others. Here we only focus on the effect this simplification has on the determinant of the Fisher matrix as this quantity measures the average parameter constraining power of the summary statistics., for all of our 21-cm summaries. It is important to note that adding this term will always increase the information content, as $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$. For more details, see the end of Appendix A.
In Fig. 7 we plot the Fisher information PDF of the commonly-used 1DPS and 2DPS. The solid curves are the same as in the previous figure, corresponding to the full expression for the Fisher information. The dashed curves only include the first term, assuming a constant covariance matrix. For both 1D and 2D power spectra, neglecting the second term underestimates the Fisher information by a factor of $\sim 2.3$ on average over the prior samples. For our 5-parameter model, this translates into an underestimate of the single parameter variance (Eq. 26) by 15%.
In Fig. 8 we compare the Fisher information PDFs with and without the second Fisher term for the remaining individual summaries. We see that for the other 21-cm summaries, the second Fisher term is even more important than for the power spectra. For example, assuming a constant covariance underestimates the mean Fisher information of the RNN for a factor of $\sim$4700 and a factor of $\sim$11100 for the wavelets, averaged over the prior samples. Note that although the IMNN is trained to normalize the covariance to unity, thus ensuring a constant covariance matrix by construction, this is only done at $\bm{\theta}_{\text{fid}}$. For the other points in the parameter space, there is no guarantee that this condition is met.
6 Conclusions
The cosmic 21-cm signal will revolutionazie our understanding of the CD and Eo). ”Big Data” analysis techniques will be required to cope with radiotelescopes such as the Square Kilometre Array (SKA). In this work we compare several approaches of compressing 21-cm lightcone data, on the basis of their Fisher information. Specifically, we compare the determinant of the Fisher information for the following summary statistics: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; (v) an information-maximizing neural network (IMNN); and (vi) the combination of 2DPS and IMNN. Importantly we compare their Fisher information across parameter space, since there is currently no standard model for the astrophysics of the first galaxies, which determine the cosmic 21-cm signal.
The 2DPS is the individual summary with the highest median Fisher information (Fig. 6, Table 1). The distribution of its Fisher information is also relatively narrow across parameter space. This means that compared to most other summaries, the 2DPS can recover astrophysical parameters with relatively similar precision regardless of what are the ”true” values of our Universe. We also show that the 2DPS slightly outperforms the 1DPS, confirming that the anisotropy of the 21-cm signal helps constrain parameters. The information content of both RNN and wavelets is generally lower than either the 1DPS or the 2DPS. However, we caution that both of these summaries could be defined in several different ways, potentially improving their performances (e.g. Greig et al. 2022, 2023; Hothi et al. 2023; Zhao et al. 2023a).
In this work, we also introduce IMNNs to the field of 21-cm cosmology. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. Indeed, we find an enormous difference between its Fisher information computed at the parameter set used for training, $\bm{\theta}_{\text{fid}}$, and its information computed at the remaining prior samples (Fig. 5). This means that the IMNN overfits to the information content of the fiducial. The regularization inherent to the IMNN loss (c.f. Eq. 22 and the discussion in Footnote 8) makes it difficult to avoid this overfitting by training over many different choices of $\bm{\theta}_{\rm fid}$. A simpler solution could be to first perform inference with a less optimal summary, in order to determine a maximum likelihood point, and then train the IMNN at that point.
Combining the IMNN with the 2DPS results in a summary whose median Fisher information is almost $\sim$4 orders of magnitude larger than the 2DPS alone. This results in a factor of $\sim$6.5–9.5 tighter single parameter variance averaged over the prior samples. This means the IMNN extracts complementary information with respect to the 2DPS. Performing simulation-based inference using such a summary could yield extremely tight and robust parameter constraints. Moreover, combining only these two complementary summaries still preserves a high level of compression, with our lightcone of $\sim 1.6\cdot 10^{6}$ cells reduced to just 77 numbers.
Finally, we stress the importance of not assuming a constant covariance when performing Fisher forecasts using 21-cm observables. This common assumption underestimates the Fisher information of the power spectrum on average by a factor of $\sim$2.3, corresponding to a 15% increase in the single parameter variance (Eq. 26). For other summaries, the assumption underestimates the Fisher information by orders of magnitude.
Acknowledgements.
D.P. thanks T. Charnock for many discussions and in depth explanations on how the IMNN function, as well as L. Makinen on help with stabilizing IMNN training. A.M. acknowledges support from the Ministry of Universities and Research (MUR) through the PRIN project ”Optimal inference from radio images of the epoch of reionization” as well as the PNRR project ”Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing”. We gratefully acknowledge computational resources of the HPC center at SNS. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support (grant IMC21cm - HP10CWIEF7). This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges-2 system, which is supported by NSF award number ACI-1928147, at the Pittsburgh Supercomputing Center (PSC).
Appendix A Proof of the determinant inequality
Let us consider two matrices $\bm{C}$ and $\bm{F}^{-1}$ of shape $m\times m$ for which the following inequality holds
$\bm{C}\geq\bm{F}^{-1}\,.$ (27)
This means that their difference is a positive, semi-definite matrix:
$\bm{x}^{T}(\bm{C}-\bm{F}^{-1})\bm{x}\geq 0\quad\forall\bm{x}\in\mathbb{R}^{m}\,.$ (28)
From this, it follows:
$\bm{x}^{T}\bm{C}\bm{x}$ $\geq\bm{x}^{T}\bm{F}^{-1}\bm{x}$ (29)
$\exp\left(-\frac{1}{2}\bm{x}^{T}\bm{C}\bm{x}\right)$ $\leq\exp\left(-\frac{1}{2}\bm{x}^{T}\bm{F}^{-1}\bm{x}\right)$ (30)
$\int\mathrm{d}\bm{x}\,\exp\left(-\frac{1}{2}\bm{x}^{T}\bm{C}\bm{x}\right)$ $\leq\int\mathrm{d}\bm{x}\,\exp\left(-\frac{1}{2}\bm{x}^{T}\bm{F}^{-1}\bm{x}\right)$ (31)
$\frac{1}{\sqrt{\det(\bm{C}/2\pi)}}$ $\leq\frac{1}{\sqrt{\det(\bm{F}^{-1}/2\pi)}}$ (32)
$\det\bm{C}$ $\geq\det\bm{F}^{-1}$ (33)
Here the trick was to push the inequality towards the multivariate Gaussian integral. After integrating, what is left is only the inverse of a normalization constant. Notice that if one wants to interpret $\bm{C}$ as a covariance matrix and $\bm{F}$ as a Fisher matrix, then it would make more sense to integrate over $\bm{x}^{T}\bm{C}^{-1}\bm{x}$ and $\bm{x}^{T}\bm{F}\bm{x}$. However this is irrelevant for the proof, as it would only result in the change of the inequality direction in each row, ending with $\det\bm{C}^{-1}\leq\det\bm{F}$. We usually assume the matrices are non-singular and thus invertible, which is needed here.
Note that the identical procedure can be used to prove that the one-term Fisher is a lower bound of the two-term Fisher. For $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ one can write $\bm{x}^{T}(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\bm{x}\geq\bm{x}^{T}\bm{F}_{\bm{\mu}}\bm{x}$, which by following the same steps translates into $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$.
Appendix B Derivation of a Gaussian Fisher information matrix
In order to derive Eq. 4, i.e. the Fisher information of a multivariate Gaussian distribution, we start by writing the Fisher information matrix equation (Eq. 1) in the equivalent form:
$\bm{F}(\bm{\theta}^{*})_{mn}=-\mathbb{E}\left[\left.\partial^{2}_{mn}\ln P(\bm{d}|\bm{\theta})\right|\bm{\theta^{*}}\right]\,,$ (34)
where we shorten the notation of derivatives as $\partial/\partial\bm{\theta}_{m}\equiv\partial_{m}$ and $\partial^{2}/\partial\bm{\theta}_{m}\partial\bm{\theta}_{n}\equiv\partial^{2}_{mn}$. This form comes from the fact that
$-\partial^{2}_{mn}\ln P(\bm{d}|\bm{\theta})=\partial_{m}\ln P(\bm{d}|\bm{\theta})\cdot\partial_{n}\ln P(\bm{d}|\bm{\theta})-\frac{1}{P(\bm{d}|\bm{\theta})}\partial^{2}_{mn}P(\bm{d}|\bm{\theta})\,.$ (35)
Taking the expectation value on both sides, together with
$\mathbb{E}\left[\left.\frac{1}{P(\bm{d}|\bm{\theta})}\partial^{2}_{mn}P(\bm{d}|\bm{\theta})\right|\bm{\theta}^{*}\right]=\partial^{2}_{mn}\int\mathrm{d}\bm{d}\,P(\bm{d}|\bm{\theta}^{*})=0\,,$ (36)
proves the equality between Eqs. 1 and 34.
Before diving into evaluating 34 for a Gaussian likelihood $\mathcal{N}(\bm{d}|\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$, we firstly list several useful relations:
$\partial_{m}\bm{\Sigma}^{-1}$ $=-\bm{\Sigma}^{-1}\,\partial_{m}\bm{\Sigma}\,\bm{\Sigma}^{-1}\,,$ (37)
$\partial_{m}|\bm{\Sigma}|$ $=|\bm{\Sigma}|\,\operatorname{tr}\left(\bm{\Sigma}^{-1}\partial_{m}\bm{\Sigma}\right)\,,$ (38)
$\mathbb{E}\left[\bm{A}(\bm{d}-\bm{\mu})\right]$ $=\bm{0}\,,$ (39)
$\mathbb{E}\left[(\bm{d}-\bm{\mu})^{T}\bm{A}(\bm{d}-\bm{\mu})\right]$ $=\operatorname{tr}(\bm{A\Sigma})\,.$ (40)
The first one follows from $\bm{\Sigma}\bm{\Sigma}^{-1}=\mathbb{1}$ and calculating its derivative, the second one is Jacobi’s relation for the derivative of the determinant. The last two are expectation relations for Gaussian distributions for an arbitrary matrix $\bm{A}$ (where $\dim\bm{A}=\dim\bm{\Sigma}$). The first of these follows from the anti-symmetry of the integral, and the second is left as an exercise for the conscientious reader.
With this at hand, we can attack the problem of calculating the Fisher matrix for a Gaussian likelihood. By taking the logarithm
$\ln P(\bm{d}|\bm{\theta})=\text{const.}-\frac{1}{2}\ln|\bm{\Sigma}|-\frac{1}{2}(\bm{d}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (41)
and computing its second derivative, we get:
$\partial^{2}_{mn}\ln P(\bm{d}|\bm{\theta})$ $=-\partial_{m}\bm{\mu}^{T}\,\bm{\Sigma}^{-1}\partial_{n}\bm{\mu}$ (42)
$\phantom{=}+\frac{1}{2}\operatorname{tr}\left(\bm{\Sigma}^{-1}\,\partial_{m}\bm{\Sigma}\,\bm{\Sigma}^{-1}\,\partial_{n}\bm{\Sigma}\right)$ (43)
$\phantom{=}-(\bm{d}-\bm{\mu})^{T}\,\bm{\Sigma}^{-1}\,\partial_{m}\bm{\Sigma}\,\bm{\Sigma}^{-1}\,\partial_{n}\bm{\Sigma}\,\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (44)
$\phantom{=}-\frac{1}{2}\operatorname{tr}\left(\bm{\Sigma}^{-1}\,\partial^{2}_{mn}\bm{\Sigma}\right)$ (45)
$\phantom{=}+\frac{1}{2}(\bm{d}-\bm{\mu})^{T}\,\bm{\Sigma}^{-1}\,\partial^{2}_{mn}\bm{\Sigma}\,\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (46)
$\phantom{=}+\partial^{2}_{mn}\bm{\mu}^{T}\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (47)
$\phantom{=}-\left(\partial_{m}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{n}\bm{\Sigma}+\partial_{n}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{m}\bm{\Sigma}\right)\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (48)
Taking the negative expectation value of the final equation can now be done by using relations 39 and 40. One can easily see that only the first three terms contribute, with others canceling each other or being equal to zero. With this, we finally have:
$\bm{F}_{mn}=\partial_{m}\bm{\mu}^{T}\,\bm{\Sigma}^{-1}\partial_{n}\bm{\mu}+\frac{1}{2}\operatorname{tr}\left(\bm{\Sigma}^{-1}\,\partial_{m}\bm{\Sigma}\,\bm{\Sigma}^{-1}\,\partial_{n}\bm{\Sigma}\right)\,.$ (49)
Appendix C Fisher information under linear transformations
Imagine we calculate the Fisher information matrix in parameter space $\bm{\theta}$ around $\bm{\theta^{*}}$:
$\bm{F}_{\bm{\theta}}(\bm{\theta^{*}})_{mn}=\mathbb{E}\left[\left.\frac{\partial}{\partial\bm{\theta}_{m}}\ln P(\bm{d}|\bm{\theta})\cdot\frac{\partial}{\partial\bm{\theta}_{n}}\ln P(\bm{d}|\bm{\theta})\right|{\bm{\theta^{*}}}\right]\,.$ (50)
If we now take a linear transformation of the parameter space $\bm{\eta}=\bm{a}\odot\bm{\theta}+\bm{b}$, where $\odot$ is element-wise multiplication, we get:
$\bm{F}_{\bm{\eta}}(\bm{\eta}^{*}=\bm{a}\odot\bm{\theta}^{*}+\bm{b})=\bm{A}\,\bm{F}_{\bm{\theta}}(\bm{\theta^{*}})\,\bm{A}\,,$ (51)
where $\bm{A}=\operatorname{diag}(\bm{a})$ is a diagonal matrix with elements on the diagonal equal to the vector $\bm{a}$. That further means:
$\det\bm{F}_{\bm{\eta}}$ $=\det{\bm{A}\bm{F}_{\bm{\theta}}\bm{A}}=(\det\bm{A})^{2}\,\det\bm{F}_{\bm{\theta}}\,,$ (52)
$\ln\det\bm{F}_{\bm{\eta}}$ $=C+\ln\det\bm{F}_{\bm{\theta}}\,.$ (53)
The final relation can be useful when thinking about Fisher information calculated in different physical units (different scalings). The only consequence of such procedure is that the logarithm of the Fisher information shifts by a constant $C$, independently on the position in the parameter space. This further means that the distribution of the Fisher information shifts by the same constant.
For the single parameter variance definition (Eq. 26), one can see the units (scaling) do not play a role, as we always express it with respect to the 2DPS, therefore:
$\frac{\sigma^{2}}{\sigma^{2}_{\rm 2DPS}}=\left(\frac{\det\bm{F}_{\bm{\theta}}}{\det\bm{F}_{\bm{\theta},{\rm 2DPS}}}\right)^{-1/5}=\left(\frac{\det\bm{F}_{\bm{\eta}}}{\det\bm{F}_{\bm{\eta},{\rm 2DPS}}}\right)^{-1/5}\,.$ (54)